OpenStudy (anonymous):

Under which condition does a country with a small GDP have a large per capita income? if it has a large population if it has a small population if the population doesn't change over time

OpenStudy (anonymous):

If a country's GDP is for example 1,000,000 with 1,000 people. the income per capita = 1,000,000/1,000 = 1,000. If the same country has suddenly 2,000 people then the income per capita becomes 1,000,000/2,000 = 500

OpenStudy (anonymous):

So, B. if it has a small population ,(The conditions that a country with a small GDP that needs to have a large per capita income, they need a small population.)